OpenStudy (anonymous):

An athlete trains for 111 min each day for as many days as possible. Write an equation that relates the number of days d that the athlete spends training when the athlete trains for 490 min.

OpenStudy (jebonna):

690/115 = d

OpenStudy (anonymous):

115*365=41975 41975/690=60.833 so he trains for 61 days

OpenStudy (anonymous):

d = 61

OpenStudy (jebonna):

d = 6 days

OpenStudy (anonymous):

jebonna is correct i read the question wrong lol sorry